OpenStudy (anonymous):

Max tossed a fair coin 3 times. What is the probability of getting heads in the first two trials and tails in the last trial?

OpenStudy (carolinar7):

1/8

OpenStudy (carolinar7):

No its not that

OpenStudy (astrophysics):

\[\text{probability} = \frac{ \text{number of favourable outcomes} }{ \text{number of possible outcomes} }\] and you should note it can never be greater than 1.

OpenStudy (astrophysics):

So if we flipped one coin for example, we would have \[\frac{ fav }{ pos } = \frac{ 1 }{ 2 } = 0.5 = 50 \%\]

OpenStudy (astrophysics):

thats the probability of getting heads if you flip one coin

OpenStudy (astrophysics):

So if you had two then we would have \[\frac{ fav }{ pos } = \frac{ 2 }{ 4 }\] see a patten here?

OpenStudy (anonymous):

so does that mean with three it would be 3/12

OpenStudy (anonymous):

oh wait maybe 3/8

OpenStudy (anonymous):

but if that is true the teacher marked it wrong

OpenStudy (irishboy123):